# Enhancing AI System Resiliency: Formulation and Guarantee for LSTM Resilience Based on Control Theory

### v4

@@ C.2.2 Data Generation
was discretized using the Runge-Kutta method with $dt=0.01$. The system parameters are shown in Table LABEL:table:system_parameters. We randomly generated a sequence of control inputs $\{u(t)\}_{t}$, set the initial state of the system to $(h_{1}(0),h_{2}(0))=(0,0)$, and obtained the system states $\{(h_{1}(t),h_{2}(t))\}_{t}$. Here, the generation of control inputs was set to a time series length of 100,000100,000100 , 000, with values randomly switched every 4,0004,0004 , 000 step following a uniform distribution $\mathcal{U}(0.5,3.0)$. Using this method, we created two independent control input sequences, one for training and one for testing.
@@
This addition of noise simulates real-world factors such as sensor measurement errors and environmental uncertainties. The noise distribution was set to $w_{1},w_{2}\sim\mathcal{N}(0,{0.1}^{2})$ for the training data and to $w_{1},w_{2}\sim\mathcal{N}(0,{0.01}^{2})$ for the test data. This is because while noise with a standard deviation of about 0.010.010.01 is expected under operational conditions, intentionally introducing larger noise during the training phase aims to improve the model’s robustness and accuracy. On the other hand, the inference target data was set using the state of the next time step as follows:
@@
The time series data are normalized to the interval [1,1][-1,1][ - 1 , 1 ] for both input $x(t)=(u(t),h_{1}(t)+w(t),h_{2}(t)+w(t))$ and output $y(t)=(h_{1}(t+1),h_{2}(t+1))$ of the LSTM before conducting inference. The equation for normalization is given below:
@@ C.2.3 Train
[…]
@@
[…]
@@ C.2.4 Evaluation
The inference accuracy was evaluated using MAE as the metric for the test data. For the time series, we calculated the MAE, ignoring the first 101010 steps and using non-normalized values for the evaluation.
Separate from the data used for evaluating inference accuracy, we generate the nominal input data $x$ and the perturbed input data $\widehat{x}$ for the estimation of recovery time $T_{R}$ using the following steps. First, we set the time series length to 350035003500 steps, the control input to $u=1.0$, and the initial water level of the system to $(h_{1}(0),h_{2}(0))=(0,0)$. Then, following the data generation process, we created $(h_{1}(t),h_{2}(t))$ and set $x(t)=(u(t),h_{1}(t),h_{2}(t))$. Then, we define $\widehat{x}$ that satisfies the assumptions of Theorem 5.3 with $t_{0}=1010$ as
@@ D.1.1 TRAIN
  • The start time is randomly selected from a uniform distribution in the range $[0,T-d]$, where $d=10$ is the duration of each pulse.
@@ D.1.2 EVALUATION
[…]
@@ D.2 Results and Analysis
Baseline DD pulse 111 DD pulse 555 DD pulse 101010 Our
@@
$\overline{T}_{R}$infty $\infty$ $\infty$ $\infty$ 19.32
@@
Table 4: Comparison of stability performance between theoretical approaches and data-driven approaches. ”DD pulse $M$” refers to a model trained with random pulses added to the training data, where $M$ indicates the upper limit of the pulse magnitude. ”Our” refers to a model trained using our proposed loss function (34). Bold numbers indicate the best performance for each metric.

### v5

@@ C.2.2 Data Generation
was discretized using the Runge-Kutta method with $dt=0.01$. The system parameters are shown in Table LABEL:table:system_parameters. We randomly generated a sequence of control inputs $\{u(t)\}_{t}$, set the initial state of the system to $(h_{1}(0),h_{2}(0))=(0,0)$, and obtained the system states $\{(h_{1}(t),h_{2}(t))\}_{t}$. Here, the generation of control inputs was set to a time series length of $100,000$, with values randomly switched every $4,000$ step following a uniform distribution $\mathcal{U}(0.5,3.0)$. Using this method, we created two independent control input sequences, one for training and one for testing.
@@
This addition of noise simulates real-world factors such as sensor measurement errors and environmental uncertainties. The noise distribution was set to $w_{1},w_{2}\sim\mathcal{N}(0,{0.1}^{2})$ for the training data and to $w_{1},w_{2}\sim\mathcal{N}(0,{0.01}^{2})$ for the test data. This is because while noise with a standard deviation of about $0.01$ is expected under operational conditions, intentionally introducing larger noise during the training phase aims to improve the model’s robustness and accuracy. On the other hand, the inference target data was set using the state of the next time step as follows:
@@
The time series data are normalized to the interval [1,1][-1,1] for both input $x(t)=(u(t),h_{1}(t)+w(t),h_{2}(t)+w(t))$ and output $y(t)=(h_{1}(t+1),h_{2}(t+1))$ of the LSTM before conducting inference. The equation for normalization is given below:
@@ C.2.3 Train
The training dataset is split into training and validation sets, with the first $50\%$ of the sequences allocated for training and the remaining $50\%$ used for validation purposes. To reduce computational load and augment data, the training dataset is created by segmenting one length time series $50,000$ into individual sequences $49,000$, each with a length of $1,000$, using a sliding window of step size $1$. On the other hand, as data augmentation is not required for the validation dataset, we create the dataset using both a window length and a step size of 1,000, solely considering computational efficiency.
@@
We used the following hyperparameters. The weights were initialized with uniform distribution $\mathcal{U}(-1/\sqrt{n_{c}},1/\sqrt{n_{c}})$ and optimized using the Adam ($\beta=(0.9,0.999)$, $\varepsilon=10^{-8}$) optimizer. The mini-batch size was $64$, the epoch size was $200$, the learning rate was $0.001$ with ReduceLROnPlateau ($\text{mode}=\text{min}$, $\text{factor}=0.1$, $\text{patience}=10$) scheduler. The baseline model was constructed with $\lambda=0$, while the stability-guaranteed model was implemented with $\lambda=1.0$. The recovery time control parameter $\varepsilon$ was selected from the set $\{0.1,0.15,0.2,0.25,0.3,0.35,0.4,0.45,0.5\}$.
@@ C.2.4 Evaluation
The inference accuracy was evaluated using MAE as the metric for the test data. For the time series, we calculated the MAE, ignoring the first $10$ steps and using non-normalized values for the evaluation.
Separate from the data used for evaluating inference accuracy, we generate the nominal input data $x$ and the perturbed input data $\widehat{x}$ for the estimation of recovery time $T_{R}$ using the following steps. First, we set the time series length to $3500$ steps, the control input to $u=1.0$, and the initial water level of the system to $(h_{1}(0),h_{2}(0))=(0,0)$. Then, following the data generation process, we created $(h_{1}(t),h_{2}(t))$ and set $x(t)=(u(t),h_{1}(t),h_{2}(t))$. Then, we define $\widehat{x}$ that satisfies the assumptions of Theorem 5.3 with $t_{0}=1010$ as
@@ D.1.1 TRAIN
  • The start time is randomly selected from a uniform distribution in the range [0,Td][0,T-d], where $d=10$ is the duration of each pulse.
@@ D.1.2 EVALUATION
where $p$ was selected from the set $\{1.0,5.0,9.0\}$, and let $\widehat{y}_{p}$ represent the output of the LSTM when $\widehat{x}_{p}$ is used as the input.
@@ D.2 Results and Analysis
Baseline DD pulse 11 DD pulse 55 DD pulse $10$ Our
@@
$\overline{T}_{R}$ $\infty$ $\infty$ $\infty$ $\infty$ 19.32
@@
Table 4: Comparison of stability performance between theoretical approaches and data-driven approaches. ”DD pulse MM” refers to a model trained with random pulses added to the training data, where $M$ indicates the upper limit of the pulse magnitude. ”Our” refers to a model trained using our proposed loss function (34). Bold numbers indicate the best performance for each metric.